Sachs Brands' defined benefit pension plan specifies annual retirement benefits equal to: 1.6% × service years × final year's salary, payable at the end of each year. Angela Davenport was hired by Sachs at the beginning of 2004 and is expected to retire at the end of 2038 after 35 years' service. Her retirement is expected to span 18 years. Davenport's salary is $94,000 at the end of 2018 and the company's actuary projects her salary to be $300,000 at retirement. The actuary's discount rate is 8%. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Required:
2.
Estimate by the projected benefits approach the amount of Davenport's annual retirement payments earned as of the end of 2018.
3. What is the company's projected benefit obligation at the end of 2018 with respect to Davenport? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)
4. If no estimates are changed in the meantime, what will be the company's projected benefit obligation at the end of 2021 (three years later) with respect to Davenport? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)Check my work 3 Sachs Brands defined benefit pension plan specifies annual retirement benefits equal to: 1.6% × service year
Answer #1

Answer:2 Calculation of the the projected benefits approach the amount of Davenport's annual retirement payments earned as of the end of 2018:

=1.6%*15 years*$94000= $22,560

Answer: 3

The present value of the retirement annuity as of the retirement date (end of 2038) is

=$22,560× 9.37189* = $211,430

* present value of an ordinary annuity of $1: n= 18, i= 8% (from PVA of $1)

It is the present value of the retirement benefits at the end of 2018:

$211,430 × 0.21455* = $45,362

* present value of $1: n= 20, i= 8% (from PV of $1)

Answer: 4

1.6 × 18 years × $94000 = $27,072

$27072 × 9.37189* = $253,716

$253716 × 0.27027** = $68572

* present value of an ordinary annuity of $1: n = 18, i = 8% (from PVA of $1)

** present value of $1: n = 17, i = 8% (from PV of $1)